# Audience Impressions of Narrative Structures and Personal Language Style in Science Communication on Social Media

### v2

@@ 6. Writer Study: Methodology
The reader study found that overall readers had no preference between narratives with and without walkthroughs but did prefer when explanations had multiple different examples to explain the topic. Previous studies have shown that scientists often struggle with two aspects of science communication on social media (1) framing science for everyday audiences and (2) using informal language to communicate science (Williams et al., 2022; Koivumäki et al., 2020). As such, we provide writers with different options to structure and frame their writing (One Example, No Example, Many Examples) and provide options to see their narrative with and without personal language to support the writing process for social media science communication. Our research questions are:
@@ 7. Results: Writer Study
7.1. RQ1: Different Narrative Structures help Writers Choose Appropriate Framing Techniques
Of the three narrative choices (One Example, Many Examples, No Examples), most science explanations used the One Example narrative to iterate on (9 of 20). The second most chosen narrative strategy was merging components between the One Example and Many Examples options (5 of 20). The Many Examples strategy was chosen 4 of 20 times. The least used structure was No Examples (2 of 20). Choices made by participants and topic are found in Table 5. Writers’ choices differed depending on the topic, their personal preferences for science communication, and whether or not they had a predefined explanation structure they wanted to follow. The One Example option often provided writers with a clear sequencing of information to follow. Some participants (P2, P4, P6, P10) chose to merge between the One Example and Many Examples and used examples covered in the Many Examples option to supplement certain explanations in the One Example narrative. Occasionally, the No Example option providing a strong technical explanation for writers to build off of.
@@ 7.1. RQ1: Different Narrative Structures help Writers Choose Appropriate Framing Techniques
Seeing different narrative structures helped 7 of 10 writers identify their own preferences for science communication by solidifying or challenging their existing beliefs (P2, P3, P5, P7, P8, P9, P10). P9 mentioned that seeing the narrative with Many Examples helped him solidify his initial feeling that “one example is a good way to engage people.” But P7 mentioned how seeing multiple structure options helped him re-evaluate his initial approach to explaining the topic of word embeddings. P7 initially wanted to use the One Example narrative structure, but instead chose to use the Many Examples narrative because the structure provided for the topic of word embeddings: “I was able to see what other things I might actually want to include when I’m trying to explain it, and what other possibilities there are for explaining.”
@@
Seeing multiple narrative structures helped 4 of 10 writers merge different framings of a topic (P2, P4, P6, P10). For multiplicative weights update, P10 merged elements from Many Examples into a base narrative with One Example to emphasize the dimensions of online learning and regret minimization which are “important facts about the topic, but wasn’t brought up in the [One Example] option.” Because the Many Examples option provided him with different angles to motivate or contextualize the topic, the selected paragraphs from Many Examples could be “slotted in at the end.”
There were 4 times where writers already had an established outline in mind for how to write the science explanation for one of their two topics (P1, P3 P8, P9). But for most writers and topics, seeing different options provided 8 of 10 writers an example of what to avoid when explaining science to an everyday audience (P1, P2, P3, P5, P6, P7, P8, P9). Seeing the No Example option helped them consider their audience’s technical ability and the background needed to explain the science. P6 said Many Examples narrative had a“rigorous depiction of what the software is, [which can be hard] for a general audience to imagine what that means.”
@@
7.2. RQ2: Different Style Options Help Writers Balance Personal Preferences with Social Media Communication
The most common narrative strategy that writers used was With Personal Language (12 of 20). Merging elements from the With Personal Language and Without Personal Language style was the second most common action that writers performed (6 of 20). Only 2 of 20 narratives used narratives without personal language. Overall, writers prefer using science narratives that include personal language and editing the LLM-generated language to match their own voice. Many writers found that comparing between narratives with and without personal language helped them identify and adopt useful techniques they might not have otherwise used in their own writing.
8 of 10 participants chose to begin with the With Personal Language option because seeing personal language provided a more accessible base narrative that writers could use to edit or replace style elements (P1, P2, P3, P5, P6, P8, P9, P10). Seeing the placement of personal language helped P6 identify where and what to edit instead of determining where and what to add if he was working from the Without Personal Language option: “I try and get a pretty good baseline, and the personal language involved adds to the baseline.” Furthermore, seeing options with personal language helped writers appreciate language that they might not have thought of using and internalize the importance of these style choices when read in juxtaposition with a narrative without personal language. For P7, seeing GPT-generated personal language helped him adapt to the genre of science communication on social media: “I think the enthusiasm, while it may not necessarily be the way that I would have written this, feels like it’s a more engaging way of trying to explain this to people on the internet.”
But for 2 of 10 writers, starting from the narrative Without Personal Language helped them evaluate the technical content of the explanation before adding back personal language (P2, P3). P2 mentioned how he prefers the remove personal language option because it is a “simpler version” that he can add his own writing voice to, instead of editing GPT’s writing voice. P3 mentioned that she chose to start with the narrative Without Personal Language because she valued the content explanation over the personal language: “I do like being more objective and professional when writing explanations about scientific concept.”
@@
Table 6. Participants, the topic they wrote on, and the corresponding narrative style that they chose. For merged narratives, “primary” denotes the narrative style that the participant chose as their base narrative, and writers incorporated elements from the remaining style option into the primary narrative.
@@ 8. Discussion
8.1. The Importance of Narratives in Science Communication
Overall, the use of an example was helpful for most readers. An example helped frame the science and provide readers a grounding for the science explanation. For certain topics, readers also enjoyed reading multiple different examples to gain an understanding of the topic through multiple perspectives. Overall the use of one or many examples was beneficial to most readers, the effectiveness of these strategies vary by reader. Some readers replace or augment science explanations with examples from their own life when the example used is unrelatable or missing. Some readers do not want an explanation at all. Different readers need different strategies to help them engage with and understand the science.
To account for varying audience preferences, we provide writers with different narrative structures to help them recognize which narrative strategy was most suitable for a given topic. Instead of having to recall science communication strategies, writers could rely on the generated structure options to identify the effective techniques that were used and merge techniques across different options. This finding connects with the design principle of recognition over recall and echoes findings from the reader study: writers benefit from seeing different narrative strategies (Du Plessis, 1994). This scaffolding technique for writers’ decision-making also exposed them to narrative structures they might not have considered and enables them to combine different structures to align better with their goals. Building off literature in constraint-based creativity, we provide writers with 3 distinct options to create a constraint space that encouraged writers to merge elements across different strategies (Tromp and Baer, 2022; Collins et al., 2024). We found that presenting different options for narrative structures helped some writers shift perspectives on how to communicate science on social media.
@@ 8.2. Presenting Personal Language as a Continuum and not a Binary
Previous research has found that scientists often feel conflicted over their personal and professional identities when engaging in science communication on social media (Koivumäki et al., 2020). Some scientists are hesitant to use colloqiual language, fearing it would undermine their authority or affect their reputation (Loroño-Leturiondo and Davies, 2018; Gero et al., 2021b). Overall, we found that readers do prefer science explanations with personal language, demonstrating that this science communication strategy on social media is helpful in engaging a broad audience. But not all readers liked or needed personal language to help them understand. As such, personal language in science communication on social media should be presented as a continuum: the effectiveness of personal language varies depending on the reader.
To help writers cater to diverse audience preferences and to accommodate writers’ own hesitancy around adopting this communication strategy, we provide writers with contrasting examples of science explanations with and without personal language. Some writers’ beliefs about how science communication should be was challenged when they viewed side-by-side explanations with and without personal language. By evaluating examples at both extremes, writers can more effectively discern which aspects of personal language align with their own communication style. Writers can easily evaluate which elements they wanted to keep, discard, or adapt to fit their own writing style. This process reduces cognitive load by offering writers with concrete reference points for how personal language can be used. Contrasting examples also allows writers to situate their own identities within a continuum, instead of trying to adhere to one method of communication. This contrast enables them to critically assess their own preferences, recognize effective techniques they may not have considered, and refine their approach based on the needs of their intended audience. These findings can help inform the design of AI-assisted writing tools to help writers explore a range of stylistic options, better understand audience preferences, and reflect on their own preferences.
@@ 9. Limitations and Future Work
Our writer study only included researchers from the computer science field. Future studies should include other STEM scientists to understand how writing in different fields of study might result in different structures and style preferences. Additionally, 9 of 10 of the participants in the writer study were men which might not offer a complete understanding of expert preferences for science communication on social media. There was little age diversity in PhD students, so additional research will need to investigate how age differences affect a scientist’s comfort with using different structure and style options for science communication on social media. When using the system, writers were not asked to actually publish their final science narratives on social media. As such, writers’ choices might not fully reflect how they would have written if they were to post on social media. Finally, future studies can also evaluate the longitudinal effects of using the system to evaluate whether the novelty of seeing different structures and style options wears off with prolonged usage and whether there are any long-term benefits in seeing multiple structure and style options during the drafting process for science communication.
@@ 10. Conclusion
It is crucial for science communication to engage the general public, and prior research suggests that using colloquial techniques from social media can be effective. Despite this, many scientists are hesitant to apply these techniques due to concerns about losing their authoritative voice. Our research highlights the complexity of public science communication and the need to balance readers’ and writers’ perspectives. While readers generally preferred explanations that included examples, walkthroughs, and personal language, their preferences were nuanced and context-dependent, influenced by their personal experiences and the complexity of the topic. Conversely, writers often feared that these techniques might compromise the clarity or authority of their explanations. However, when given the opportunity to explore various narrative structures and styles, writers were able to navigate their choices with greater confidence, finding a balance between colloquial and formal approaches. This suggests that effective science communication benefits from exploring diverse options, allowing writers to tailor their style to the scientific topic, their own preferences, and the needs of their audience.

### v3

@@ 6. Writer Study: Methodology
From the reader study, we found that readers had personal preferences for how different public science communication strategies (example, walkthrough, and personal language) were used. Readers said that examples and walkthroughs helped provide more a scaffolded understanding of the STEM topics, but not all topics or readers needed the same level of elaboration. For personal language, while many readers preferred it, there was a spectrum of preferences regarding how much personal language was needed in science narratives.
The findings from the reader study illustrate that science communication strategies cannot be applied as one-size-fits-all approach: use or not use examples/walkthroughs/personal language. Instead, the effectiveness of these strategies depends on both the STEM topic and the individual reader’s preferences. Given the diverse and contextual preferences that readers had for these science communication strategies, static writing guidelines do not accommodate these preferences. As such, we designed a public science writing support tool that assists scientists in negotiating how to adopt science communication strategies, with the goal of helping scientists make more informed and intentional choices about how to communicate STEM topics.
In designing a system for writers, we use a two step process that helps writers first structure, then style their science narrative. Structure pertains to how the science is communicated through examples and walkthroughs. Style refers to the tone and voice of the narrative. Our system uses contrastive examples to help writers navigate structure (one, none, and many examples) and style (with and without personal language) options. We conduct a user study with 10 scientists to understand how seeing different examples might help writers by providing them with contrastive versions of narratives that vary in structure (e.g., examples and walkthroughs) and style (e.g., personal language).
Specifically we are interested in the following research questions:
@@ 7. Results: Writer Study
7.1. RQ1: Different Narrative Structures Helps Scientists Identify What Content to Include (or Not Include)
Scientists chose different narrative structures to work off based on the STEM topic they were writing about, their individual preferences, or their judgments about what would be most accessible to an everyday audience. Most science narratives used the One Example narrative structure to iterate on (13 of 20). The Many Examples strategy was chosen 5 of 20 times. The least used structure was No Examples (2 of 20). Choices made by participants and topic are found in Table 5. From the interviews, scientists’ choice of narrative structure depended on the topic and the content that they wanted to include in their science narrative. Scientists chose the One Example option because the structure provided a clear sequence of explanation for scientists follow. Other scientists chose to merge between the One Example and Many Examples structure to provide additional examples to explain certain aspects of a STEM topic. One writer chose to use the No Example option because it provided a strong technical explanation for scientists to build off of (P3).
@@
Scientists’ preferences and beliefs about how to structure science narratives were revealed or challenged when they saw contrasting structure options (P2, P3, P5, P7, P8, P9, P10). Seeing the comparison of different narrative structures, helped P9 solidify his initial feeling that “one example is a good way to engage people.” In contrast, P7’s initial inclination to use the One Example narrative structure was challenged after seeing the different structure options. P7 chose to use the Many Examples narrative structure because he was “able to see what other things I might actually want to include when I’m trying to explain it, and what other possibilities there are for explaining.” Seeing multiple different narrative structures prompted P6 to more carefully evaluate which structure option was most suitable for his topic of formal verification:“I’m not sure if I will pick the one example narrative, because it’s a hard topic to explain to a general audience. I’ll look through the other options more to see if there’s something more suitable for this topic.” Contrasting narrative structures helped confirm existing writer preferences and prompted some scientists to reconsider which narrative structure was most suitable for their topic.
The narrative structures that scientists did not utilize often provided a negative example that highlighted what to avoid when explaining science to an everyday audience for 8 scientists (P1, P2, P3, P5, P6, P7, P8, P9). P6 said Many Examples narrative had a ‘‘rigorous depiction of what the software is, [which can be hard] for a general audience to imagine what that means.” For P9, seeing the No Example narrative helped him think from the perspective of a reader and consider “why [a reader] would decide to click on the [No Example narrative]. Unless the reader was looking to understand what merge sort was, the [No Example narrative] would not help the everyday reader get into the topic.” Even though scientists may not have chosen a particular narrative structure to use, the act of reading through different narrative options helped scientists clarify how to present their science topic for an everyday audience.
@@
Four scientists chose to merge two narrative structures together to combine the different strengths of each structure (P2, P4, P6, P10). For the topic of multiplicative weights update, P10 merged elements from Many Examples into a base narrative with One Example to emphasize the dimensions of online learning and regret minimization which are “important facts about the topic, but wasn’t brought up in the [One Example] option.” Because the Many Examples option provided him with different angles to motivate or contextualize the topic, the selected paragraphs from Many Examples could be “slotted in at the end.” Seeing different narrative structures helped scientists balance trade-offs such as technical content and clarity.
Seeing different structure examples helped scientists identify the science content that they might want to include and provide a scaffold to easily integrate different content that they should include for the STEM topic they were writing about.
@@
7.2. RQ2: Different Style Options Help Scientists Balance Personal Preferences with Social Media Communication
The most common narrative strategy that scientists used was With Personal Language (12 of 20). Merging elements from the With Personal Language and Without Personal Language style was the second most common action that scientists performed (6 of 20). Only 2 of 20 narratives used narratives without personal language. From interviews, scientists generally preferred using science narratives with personal language and editing the LLM-generated language to match their own voice. Many scientists found that comparing between narratives with and without personal language helped them identify and adopt useful techniques they might not have otherwise used in their own writing. Even though most scientists choose to iterate on science narrative with personal language, some scientists found that working off of narratives without personal language helped them clarify the science before adding back personal language.
Scientists’ recognized the importance of personal language and integrated personal language into their science narrative after seeing narratives with and without personal language. Even when scientists were initially hesitant about using personal language, seeing narratives with and without personal language helped four scientists adopt personal language in their final narratives (P1, P3, P9, P10). For P7, seeing GPT-generated personal language helped him adapt to the genre of science communication on social media: “I think the enthusiasm, while it may not necessarily be the way that I would have written this, feels like it’s a more engaging way of trying to explain this to people on the internet.” P1 was initially resistant to using personal language, but after reading the “Without Personal Language” narrative, P1 said “the comparison between with and without personal language, was good as a recall to me that this type of personal language is important in a twitter threads.” Contrasting examples with and without personal language helped scientists recognize and adopt personal language, even when they were initially hesitant to including it.
Furthermore, seeing the two contrasting style options helped 7 scientists balance technical content with personal exposition (P1, P2, P3, P5, P7, P8, P9). Narratives with personal language often contained more personal exposition about specific details or setting the scene, so seeing a narrative without personal language helped scientists “introduce the premise a lot more concisely” (P7). P9 said that “because the narrative without personal language is more technical as I’m editing the narrative with personal language, I’ll go over to the right [without personal language] if I feel like more of an explanation is needed, and see if there’s a technical detail that I can bring in.” The tradeoff between personal language and concision was highlighted through the side-by-side comparison of science narratives with and without personal language.
Seeing contrasting examples of science narratives with and without personal language helped scientists (1) adopt the use of personal language or (2) strike a balance between technical concision and descriptive problem statements.
@@
Table 6. Participants, the topic they wrote on, and the corresponding narrative style that they chose. For merged narratives, “primary” denotes the narrative style that the participant chose as their base narrative, and scientists incorporated elements from the remaining style option into the primary narrative.
@@ 8. Discussion
8.1. Narrative Structure Options Help Scientists Identify Science Content to Include
Our findings demonstrate that exposing scientists to multiple narrative structures plays a role in shaping what content scientists choose to include or exclude. Since not all STEM topics require the same level of detail, providing multiple different structure options allows scientists to identify the level of specificity that is important for their STEM topic. Seeing different narrative structure examples helps scientists identify gaps in the explanation to add more detail—by including alternative examples to explain the topic—or to remove extraneous details to support readers in understanding the science. Furthermore, seeing different narrative structures puts the writer into the perspective of a reader and helps them consider what types of science content would be the most beneficial to an everyday audience. For scientists who have already amassed technical expertise in their domain, using AI to present initial drafts allows them to focus on evaluating the trade-offs between different structure options such as technical detail and clarity.
This finding relates to a broader usability heuristic of “recognize over recall” which emphasizes the use of visual cues to help with memory retrieval as opposed to remembering something from scratch (Gigerenzer and Goldstein, 2011; Nielsen, 10). This design heuristic is important because it reduces the amount of cognitive effort required by the user to perform the task. In the context of science communication, presenting scientists with explicit structural alternatives acts as a recognition aid: instead of having to recall different ways of structuring an explanation, scientists can recognize, compare, and adapt pre-existing narrative structure options. This not only lowers the effort required for scientists to use effective communication practices but also helps scientists to critically evaluate the tradeoffs between different structure options to find one (or a combination of structures) to best communicate their topic to an everyday audience.
@@ 8.2. Presenting Personal Language as a Continuum and not a Binary
Prior research has shown that telling scientists to use personal language is not enough to motivate them to adopt this science communication strategy (Gero et al., 2021b). We found that presenting starkly contrasting narratives with and without personal language helped scientists identify how much personal language they wanted to include, even though it was typically in between the examples. Even scientists who were initially hesitant to use personal language in their science narrative adopted some personal language in their final science narrative because they recognized how narratives without personal language were overly technical and not engaging. Contrasting examples allowed scientists to situate their own identities within a continuum, recognize effective techniques they may not have considered, and refine their approach based on the needs of their intended audience. This highlights the need to balance between different writing goals: too little personal language may distance readers, while too much may overshadow the science, and there is no single “right” amount that universally excites or engages audiences.
Exposing scientists to two extremes (with and without personal language) served as anchor points that enabled scientists to situate their own style choices in relation to two extreme science communication styles. This suggests that effective science communication strategies should not prescribe personal language as an all-or-nothing approach, but instead provide opportunities for scientists to choose the degree of personal language that best matches their topic, audience, and communication goals. More broadly, these findings point to the value of presenting writing guidelines as a continuum, where contrastive options can help writers reflect on and adopt strategies in ways that align with their own voice and writing context. AI-assisted writing tools can embed this principle by surfacing stylistic contrasts that guide, rather than prescribe, how these communication guidelines can be implemented in practice.
@@ 9. Limitations and Future Work
Our writer study only included researchers from the computer science field. Future studies should include other STEM scientists to understand how writing in different fields of study might result in different structures and style preferences. Additionally, 9 of 10 of the participants in the writer study were men which might not offer a complete understanding of expert preferences for science communication on social media. There was little age diversity in PhD students, so additional research will need to investigate how age differences affect a scientist’s comfort with using different structure and style options for science communication on social media. When using the system, scientists were not asked to actually publish their final science narratives on social media. As such, scientists’ choices might not fully reflect how they would have written if they were to post on social media. Finally, future studies can also evaluate the longitudinal effects of using the system to evaluate whether the novelty of seeing different structures and style options wears off with prolonged usage and whether there are any long-term benefits in seeing multiple structure and style options during the drafting process for science communication.
@@ 10. Conclusion
It is crucial for science communication to engage the general public, and prior research suggests that using colloquial techniques from social media can be effective. Despite this, many scientists are hesitant to apply these techniques due to concerns about losing their authoritative voice. Our research highlights the complexity of public science communication and the need to balance readers’ and scientists’ perspectives. While readers generally preferred explanations that included examples, walkthroughs, and personal language, their preferences were nuanced and context-dependent, influenced by their personal experiences and the complexity of the topic. Conversely, scientists often resisted adopting these techniques, worried that these strategies might compromise the clarity or authority of their explanations. However, when given the opportunity to explore various narrative structure and style options, scientists were able to reflect on trade-offs, merge across options, and situate themselves along a continuum rather than treating strategies as rigid rules. These findings suggest that effective science communication support should not prescribe fixed guidelines, but instead present contrastive options that help scientists balance engagement and clarity, authority and relatability.